# Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases

### v2

@@ Abstract
Topological phase transitions beyond anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) and double semion (DS) phases, which has two distinct $\mathbb{Z}_{2}$ topological orders in (2+1)(2+1)D. Previous studies reveal that the transition between them can be either first order or via an intermediate phase, thus the existence of a directly continuous transition between them remains a long-standing problem. Motivated by the fact that both phases can arise from condensing distinct anyons in the $\mathbb{Z}_{4}$ topological order, we introduce a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model to study the TC-DS transition. We confirm the existence of a continuous (2+1)(2+1)D XY* transition between the TC and DS phases by mapping it to a two-coupled quantum Ising model. Importantly, using the condensation order parameters and the area law coefficients of the Wilson loops, we further reveal that $\mathbb{Z}_{4}$ anyons, fractionalized from the $\mathbb{Z}_{2}$ topological orders, become deconfined at the transition between $\mathbb{Z}_{2}$ topologically ordered phases. Our results open a path toward developing a theoretical framework for topological phase transitions beyond anyon condensation.
Introduction. Topological phases of matter, first discovered in the fractional quantum Hall effect [1, 2], have revolutionized our understanding of quantum many-body systems. Owing to their potential applications in quantum error correction and fault-tolerant quantum computing [3, 4], topological phases have become a central focus of both theoretical and experimental research. Exactly solvable models with topologically ordered ground states have been constructed to study these phases [4, 5, 6], and some have been experimentally realized using quantum processors and quantum simulators [7, 8, 9, 10]. As phases of matter beyond the conventional Ginzburg-Landau-Wilson (GLW) paradigm of spontaneous symmetry breaking, topologically ordered states pose significant challenges for studying their phase transitions within the framework of the GLW paradigm. Nonetheless, topological phase transitions driven by anyon condensation can be partially understood using the GLW paradigm, because they can be mapped, via duality transformations, to spontaneous symmetry-breaking phase transitions characterized by local order parameters [11, 12, 13, 14, 15, 16], and such topological phase transitions have been extensively studied using various models and methods [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 14, 15, 27, 28].
@@
However, topological phase transitions that are not induced by anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) [4] and double semion (DS) phases [6], which represent two distinct $\mathbb{Z}_{2}$ topological orders with incompatible anyons. By interpolating fixed TC and DS string-net models using a one-parameter Hamiltonian, the transition has been studied using exact diagonalization [29] and quantum Monte Carlo [30, 31], the former method indicates a first-order transition, while the latter method suggests the existence of an intermediate gapless phase, as shown in Fig. 1a. Additionally, a fine-tuned transition between the TC and DS phases can be realized using deformed projected entangled pair states (PEPS) with some tuning parameters [32, 33, 34, 35]. Furthermore, a field-theoretic argument has been proposed for a continuous transition between the TC and DS phases in the presence of SU(2) symmetry [36]. Despite these efforts, how to concretely realize a direct continuous transition between the TC and DS phases using a generic Hamiltonian remains a long-standing open question. If such a continuous transition exists, the nature of the critical point, particularly how the incompatible anyons of the two $\mathbb{Z}_{2}$ topological orders reconcile at the critical point, is unclear.
@@
In this letter, we confirm the existence of a direct continuous TC-DS transition and provide new insights into its nature. Rather than considering a model interpolating directly between fixed points of the TC and DS phases [29, 30, 31], we propose a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model based on the observation that both phases can emerge from the $\mathbb{Z}_{4}$ topological order via anyon condensation [20, 22, 37]. To determine its phase diagram and the quantum criticality of the continuous TC-DS transition, we introduce a duality transformation that maps the perturbed $\mathbb{Z}_{4}$ QD model to a two-coupled (2+1)(2+1)D quantum Ising model and approximate its ground states with variationally optimized infinite PEPS (iPEPS) [38, 39] based on automatic differentiation [40].
@@
Using the condensation order parameters and the area law coefficients of the Wilson loop expectation values, we identify that the continuous TC-DS transition belongs to the XY* universality class. Interestingly, as shown in Fig. 1b, the continuous TC-DS transition resembles the deconfined quantum critical point [41, 42]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, the perturbed $\mathbb{Z}_{4}$ QD model has two different 1-form symmetries, which are loop-like generalized symmetries. One 1-form symmetry is broken spontaneously in the TC phase, while the other one is broken spontaneously in the DS phase. Second, some anyons in the $\mathbb{Z}_{2}$ topological orders are fractionalized and become deconfined at the continuous TC-DS transition point. Our results suggest that at the critical point between topological phases that cannot be connected via anyon condensation, fractionalized anyons that are confined in both adjacent phases can emerge. This opens a path toward developing a general theoretical framework for topological phase transitions beyond anyon condensation.
@@
A perturbed $\mathbb{Z}_{4}$ quantum double model. Because the model interpolating between the fixed-point Hamiltonians of the TC and DS phases [29, 30] is difficult to analyze analytically and simulate numerically, we alternatively propose a perturbed $\mathbb{Z}_{4}$ QD model to realize the topological phase transition between TC and DS phases  [^1] The $\mathbb{Z}_{4}$ QD model is also known called $\mathbb{Z}_{4}$ toric code model. In this work, “TC” refers exclusively to the $\mathbb{Z}_{2}$ toric code model, and we avoid calling $\mathcal{H}_{\text{QD}}$ the $\mathbb{Z}_{4}$ TC model to prevent confusion.. The fixed-point Hamiltonian for the $\mathbb{Z}_{4}$ QD phase on the square lattice is [4]:
@@
The Hamiltonian possesses an extensive set of conserved quantities: $A^{2}_{v},H_{\text{QD}}]=0$ for all vertices, and $B^{2}_{p},H_{\text{QD}}]=0$ for all plaquettes. Multiplying these operators generates the so-called 1-form symmetries [52, 53, 54], represented by the non-local loop operators $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}W_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, where $C$ and $\hat{C}$ denote loops on the original and dual lattices, respectively. Since different topological orders can be understood as distinct patterns of higher-form symmetry breaking [55], the 1-form symmetries can be used to characterize different topological phases.
@@
Condensation order parameters and area law coefficients. We can use non-local order parameters to detect spontaneous 1-form symmetry breaking, which implies different topological phases. A kind of such non-local order parameters includes $\left\langle\prod_{e\in S}Z_{e}^{2}\right\rangle$, $\left\langle\prod_{e\in\hat{S}}W_{e}^{2}\right\rangle$, and $\left\langle\prod_{e\in\hat{S}}X_{e}^{2}\right\rangle$ [56, 57], where $S$ ($\hat{S}$) denotes the open strings on the original (dual) lattices. They can be viewed as segments of closed loops $C$ ($\hat{C}$), as detailed in SM [44]. Because the anyons $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ are created at the endpoints of the string operators, and their non-zero expectation values imply the corresponding anyons condense, as summarized in Tab. 1, and we call them condensation order parameters.
@@
Cond. $\langle\pmb{e}^{2}\rangle$ $\left|\left\langle\prod_{e\in S}Z^{2}_{e}\right\rangle\right|$ $\left|\left\langle\tau_{i}^{z}\right\rangle\right|$ 0 0 >0>0 >0>0
$\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}W^{2}_{e}\right\rangle\right|$ $\left|\left\langle\tau_{i}^{z}\sigma_{i}^{x}\right\rangle\right|$ 0 >0>0 0 >0>0
$\langle\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}X^{2}_{e}\right\rangle\right|$ $\left|\left\langle\sigma_{i}^{x}\right\rangle\right|$ 0 0 0 >0>0
Conf. $\alpha(\pmb{e})$ $\frac{\log\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle}{-a_{C}}$ $\frac{\log\langle\prod_{i\in R}\sigma_{i}^{z}\rangle}{-a_{R}}$ 0 >0>0 0 >0>0
$\alpha(\pmb{e}\pmb{m})$ $\frac{\log\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle}{-a_{\hat{C}}}$ $\frac{\log\langle\prod_{i\in R}\tau^{x}_{i}\sigma^{z}_{i}\rangle}{-a_{R}}$ 0 0 >0>0 >0>0
$\alpha(\pmb{m})$ $\frac{\log\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle}{-a_{\hat{C}}}$ $\frac{\log\langle\prod_{i\in R}\tau^{x}_{i}\rangle}{-a_{R}}$ 0 >0>0 >0>0 >0>0
Another kind of non-local order parameters is defined using Wilson loop operators $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}\right\rangle$, where $e^{\prime}$ denotes the edge immediately southwest of $e$, and the exponents $\gamma(e),\leavevmode\nobreak\ \hat{\gamma}(e)$ taking values $1$ or $3$ depend on the edge orientations [37, 44]. It is well known that the expectation values of the Wilson loop operators can detect the confinement of anyons [58, 59]. Let us focus on $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, where the loop operator can be regarded as creating the pair $\pmb{e}$ and $\pmb{e}^{3}$ and annihilating them. When the expectation value satisfies $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle\propto\mathrm{e}^{-\alpha(\pmb{e})a_{C}}$ for a large loop $C$ that surrounds the area $a_{C}$, where $\alpha(\pmb{e})$ is the area law coefficient. When $\alpha(\pmb{e})>0$, it is the area law and it indicates $\pmb{e}$ is confined. If $\alpha(\pmb{e})=0$, it implies that $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$ decays at most as the perimeter of the loop $C$, i.e. the perimeter law, indicating that $\alpha(\pmb{e})$ is deconfined. Therefore, we can use the area law coefficient $\alpha(\pmb{e})$ to distinguish the deconfinement and confinement of $\pmb{e}$. A similar argument applies to $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle$, which is summarized in Tab. 1.
@@
see detailed derivation in SM [44]. If the term $\sum_{i}\tau_{i}^{x}\sigma^{z}_{i}$ is absent and $h_{x}=h_{z}$, it becomes a standard (2+1)(2+1)D quantum Ashkin-Teller model. It is obvious that the phase diagram of $H_{\text{dual}}$ is symmetric about $h_{x}=h_{z}$, because the model is invariant under the exchange of $\sigma_{i}^{x}\leftrightarrow\tau_{i}^{z}$, $\sigma_{i}^{z}\leftrightarrow\tau_{i}^{x}$, and $h_{x}\leftrightarrow h_{z}$. Furthermore, its phase diagram is also symmetric about $h_{z}=h_{w}$, because $UH_{\text{dual}}(h_{x},h_{z},h_{w})U^{\dagger}=H_{\text{dual}}(h_{x},h_{w},h_{z})$, where the unitary transformation is $U=\prod_{i}\left(\frac{1+\sigma_{i}^{x}}{2}+\tau_{i}^{x}\frac{1-\sigma_{i}^{x}}{2}\right)$. Since duality transformations preserve the structure of phase diagrams, we can directly deduce the phase diagram of the $\mathbb{Z}_{4}$ QD model $H_{\text{QD}}$ from that of $H_{\text{dual}}$.
The non-local condensation order parameters and the Wilson loop expectation values of the perturbed $\mathbb{Z}_{4}$ QD model can be mapped to the local order parameters and the membrane-like disorder parameters of the two-coupled quantum Ising model [61, 62] (see SM [44] and Tab. 1). This allows us to extract condensation order parameters and the area law coefficients directly from the ground states of the two-coupled Ising model, which can be approximated using $C_{4v}$-symmetric iPEPS [39, 38, 40]. Compared to other methods such as quantum Monte Carlo, iPEPS is more suitable to evaluate non-local order parameters, and the method of extracting area law coefficients from iPEPS is proposed in Ref. [63]. By comparing the numerical results of the condensation order parameters and the area law coefficients with Tab. 1, we obtain the phase diagrams for various $h_{x}$ planes (see SM [44]), one of which is shown in Fig. 3. It includes the $\mathbb{Z}_{4}$ QD, TC, DS, and trivial phases, and their phase boundaries are described by the 3D Ising* universality class describing 1-form symmetry breaking transitions [64]. Next, we focus on the multi-critical point, especially the continuous TC–DS transition.
@@
Fig. 4a shows the area law coefficients $\alpha(\pmb{e})$, $\alpha(\pmb{e}\pmb{m})$ and $\alpha(\pmb{m})$ are zero when $h_{w}=h_{z}<h_{c}\approx 1.490$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{n}$, $\pmb{e}^{n}\pmb{m}^{k}$ and $\pmb{m}^{k}$ ($n,k\in\{1,3\}$) are confined. Moreover, Fig. 4b shows the condensation order parameters $\langle\pmb{e}^{2}\rangle$, $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{m}^{2}\rangle$ are zero when $h_{w}=h_{z}<h_{c}$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ condense. By comparing the numerical results in Figs. 4a and b with Tab. 1, we conclude the perturbed $\mathbb{Z}_{4}$ QD model is in the $\mathbb{Z}_{4}$ QD when $h_{w}=h_{z}<h_{c}$ and trivial phases when $h_{w}=h_{z}>h_{c}$.

### v3

@@ Abstract
Topological phase transitions beyond anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) and double semion (DS) phases, which has two distinct $\mathbb{Z}_{2}$ topological orders in (2+1)(2+1)D. Previous studies reveal that the transition between them can be either first order or via an intermediate phase, thus the existence of a direct continuous transition between them remains a long-standing problem. Motivated by the fact that both phases can arise from condensing distinct anyons in the $\mathbb{Z}_{4}$ topological order, we introduce a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model to study the TC-DS transition. We confirm the existence of a continuous (2+1)(2+1)D XY* transition between the TC and DS phases by mapping it to a two-coupled quantum Ising model. Importantly, using the condensation order parameters and the area law coefficients of the Wilson loops, we further reveal that $\mathbb{Z}_{4}$ anyons, fractionalized from the $\mathbb{Z}_{2}$ topological orders, become deconfined at the transition between $\mathbb{Z}_{2}$ topologically ordered phases. Our results open a path toward developing a theoretical framework for topological phase transitions beyond anyon condensation.
Introduction. Topological phases of matter, first discovered in the fractional quantum Hall effect [1, 2], have revolutionized our understanding of quantum many-body systems. Owing to their potential applications in quantum error correction and fault-tolerant quantum computing [3, 4], topological phases have become a central focus of both theoretical and experimental research. Exactly solvable models with topologically ordered ground states have been constructed to study these phases [4, 5, 6], and some of them have been experimentally realized using quantum processors and quantum simulators [7, 8, 9, 10]. As phases of matter beyond the conventional Ginzburg-Landau-Wilson (GLW) paradigm of spontaneous symmetry breaking, topologically ordered states pose significant challenges for studying their phase transitions within the framework of the GLW paradigm. Nonetheless, topological phase transitions driven by anyon condensation can be partially understood using the GLW paradigm, because they can be mapped, via duality transformations, to spontaneous symmetry-breaking phase transitions characterized by local order parameters [11, 12, 13, 14, 15, 16]. Such topological phase transitions have been extensively studied using various models and methods [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 14, 15, 27, 28].
@@
However, topological phase transitions that are not induced by anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) [4] and double semion (DS) phases [6], which represent two distinct $\mathbb{Z}_{2}$ topological orders with incompatible anyons. By interpolating the fixed-point TC and DS string-net models using a one-parameter Hamiltonian, the transition has been studied using exact diagonalization [29] and quantum Monte Carlo [30, 31], the former method indicates a first-order transition, while the latter method suggests the existence of an intermediate gapless phase, as shown in Fig. 1a. Additionally, a fine-tuned transition between the TC and DS phases can be realized using deformed projected entangled pair states (PEPS) with suitable tuning parameters [20, 22, 32, 33]. Furthermore, a field-theoretic argument has been proposed for a continuous transition between the TC and DS phases in the presence of SU(2) symmetry [34]. Despite these efforts, how to concretely realize a direct continuous transition between the TC and DS phases using a generic Hamiltonian remains a long-standing open question. If such a continuous transition exists, the nature of the critical point, particularly how the incompatible anyons of the two $\mathbb{Z}_{2}$ topological orders reconcile at the critical point, remains unclear.
@@
In this letter, we confirm the existence of a direct continuous TC-DS transition and provide new insights into its nature. Rather than considering a model interpolating directly between fixed points of the TC and DS phases [29, 30, 31], we propose a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model based on the observation that both phases can emerge from the $\mathbb{Z}_{4}$ topological order via anyon condensation [20, 22, 35]. To determine its phase diagram and the quantum criticality of the continuous TC-DS transition, we introduce a duality transformation that maps the perturbed $\mathbb{Z}_{4}$ QD model to a two-coupled (2+1)(2+1)D quantum Ising model, and approximate its ground states using variationally optimized infinite PEPS (iPEPS) [36, 37] with automatic differentiation [38].
@@
Using the condensation order parameters and the area law coefficients of the Wilson loop expectation values, we identify that the continuous TC-DS transition belongs to the XY* universality class. Interestingly, as shown in Fig. 1b, the continuous TC-DS transition resembles a deconfined quantum critical point [39, 40]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, the perturbed $\mathbb{Z}_{4}$ QD model has two distinct 1-form symmetries, which are loop-like generalized symmetries. One 1-form symmetry is broken spontaneously in the TC phase, while the other one is broken spontaneously in the DS phase. Second, some anyons in the $\mathbb{Z}_{2}$ topological orders are fractionalized and become deconfined at the continuous TC-DS transition point. Our results suggest that, at the critical point between topological phases that cannot be connected via anyon condensation, fractionalized anyons that are confined in both adjacent phases can emerge. This opens a path toward developing a general theoretical framework for topological phase transitions beyond anyon condensation.
@@
A perturbed $\mathbb{Z}_{4}$ quantum double model. Because the model interpolating between the fixed-point Hamiltonians of the TC and DS phases [29, 30] is difficult to analyze analytically and simulate numerically, we alternatively propose a perturbed $\mathbb{Z}_{4}$ QD model to realize the topological phase transition between the TC and DS phases  [^1] The $\mathbb{Z}_{4}$ QD model is also known called $\mathbb{Z}_{4}$ toric code model. In this work, “TC” refers exclusively to the $\mathbb{Z}_{2}$ toric code model, and we avoid calling $\mathcal{H}_{\text{QD}}$ the $\mathbb{Z}_{4}$ TC model to prevent confusion.. The fixed-point Hamiltonian for the $\mathbb{Z}_{4}$ QD phase on the square lattice is [4]:
@@
The Hamiltonian possesses an extensive set of conserved quantities: $A^{2}_{v},H_{\text{QD}}]=0$ for all vertices, and $B^{2}_{p},H_{\text{QD}}]=0$ for all plaquettes. Products of these operators generate the so-called 1-form symmetries [50, 51, 52], represented by the non-local loop operators $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}W_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, where $C$ and $\hat{C}$ denote loops on the original and dual lattices, respectively. Since different topological orders can be understood as distinct patterns of higher-form symmetry breaking [53], the 1-form symmetries can be used to characterize different topological phases.
@@
Condensation order parameters and area law coefficients. We use non-local order parameters to detect spontaneous 1-form symmetry breaking, which distinguishes different topological phases. One class of such non-local order parameters includes $\left\langle\prod_{e\in S}Z_{e}^{2}\right\rangle$, $\left\langle\prod_{e\in\hat{S}}W_{e}^{2}\right\rangle$, and $\left\langle\prod_{e\in\hat{S}}X_{e}^{2}\right\rangle$ [54, 55], where $S$ ($\hat{S}$) denotes the open strings on the original (dual) lattices. They can be viewed as segments of closed loops $C$ ($\hat{C}$), as detailed in SM [42]. Because the anyons $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ are created at the endpoints of the string operators, their non-zero expectation values imply the corresponding anyons condense, as summarized in Tab. 1. We therefore call them condensation order parameters.
@@
Cond. $\langle\pmb{e}^{2}\rangle$ $\left|\left\langle\prod_{e\in S}Z^{2}_{e}\right\rangle\right|^{1/2}$ $\left|\left\langle\tau_{i}^{z}\right\rangle\right|$ 0 0 >0>0 >0>0
$\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}W^{2}_{e}\right\rangle\right|^{1/2}$ $\left|\left\langle\tau_{i}^{z}\sigma_{i}^{x}\right\rangle\right|$ 0 >0>0 0 >0>0
$\langle\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}X^{2}_{e}\right\rangle\right|^{1/2}$ $\left|\left\langle\sigma_{i}^{x}\right\rangle\right|$ 0 0 0 >0>0
Conf. $\alpha(\pmb{e})$ $\frac{\ln\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle}{-a_{C}}$ $\frac{\ln\langle\prod_{i\in R}\sigma_{i}^{z}\rangle}{-a_{R}}$ 0 >0>0 0 >0>0
$\alpha(\pmb{e}\pmb{m})$ $\frac{\ln\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle}{-a_{\hat{C}}}$ $\frac{\ln\langle\prod_{i\in R}\tau^{x}_{i}\sigma^{z}_{i}\rangle}{-a_{R}}$ 0 0 >0>0 >0>0
$\alpha(\pmb{m})$ $\frac{\ln\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle}{-a_{\hat{C}}}$ $\frac{\ln\langle\prod_{i\in R}\tau^{x}_{i}\rangle}{-a_{R}}$ 0 >0>0 >0>0 >0>0
Another class of non-local order parameters is defined using Wilson loop operators $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}\right\rangle$, where $e^{\prime}$ denotes the edge immediately southwest of $e$, and the exponents $\gamma(e),\penalty 10000\ \hat{\gamma}(e)$, taking values $1$ or $3$, depend on the edge orientations [35, 42]. It is well known that the expectation values of the Wilson loop operators can detect the confinement of anyons [15, 57]. Let us focus on $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, where the loop operator creates a pair of anyons $\pmb{e}$ and $\pmb{e}^{3}$ and then annihilates them. The expectation value satisfies $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle\propto\mathrm{e}^{-\alpha(\pmb{e})a_{C}}$ for a large loop $C$ that surrounds the area $a_{C}$, where $\alpha(\pmb{e})$ is the area law coefficient. When $\alpha(\pmb{e})>0$, it is the area law, indicating that $\pmb{e}$ is confined. If $\alpha(\pmb{e})=0$, it implies that $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$ decays at most as the perimeter of the loop $C$, i.e. the perimeter law, indicating that $\pmb{e}$ is deconfined. Therefore, we can use the area law coefficient $\alpha(\pmb{e})$ to distinguish the deconfinement and confinement of $\pmb{e}$. A similar argument applies to $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle$, which is summarized in Tab. 1.
@@
When taking the boundary condition into account, i.e., consider the system on a torus, the duality mapping becomes $H_{\text{QD}}=\bigoplus_{b^{\sigma}_{x},b^{\sigma}_{y},b^{\tau}_{x},b^{\tau}_{y}}\frac{1+\prod_{i}\sigma^{z}_{i}}{2}\frac{1+\prod_{i}\tau^{x}_{i}}{2}H_{\text{dual}}(b^{\sigma}_{x},b^{\sigma}_{y},b^{\tau}_{x},b^{\tau}_{y}),$ where $b^{\sigma}_{x},b^{\sigma}_{y},b^{\tau}_{x},b^{\tau}_{y}$ are boundary conditions in $x$ and $y$ directions of the $\sigma$ or $\tau$ spins which can be either periodic (++) or antiperiodic (-), and the anti-periodic boundary condition means that the ferromagnetic couplings along some non-contractible loops are changed to the anti-ferromagnetic couplings. In the SM [42], we further analyze that the symmetry sectors and boundary conditions induced by the duality transformation do not affect the evaluation of order parameters. If the term $\sum_{i}\tau_{i}^{x}\sigma^{z}_{i}$ is absent and $h_{x}=h_{z}$, it becomes a standard (2+1)(2+1)D quantum Ashkin-Teller model. It is obvious that the phase diagram of $H_{\text{dual}}$ is symmetric about $h_{x}=h_{z}$, because the model is invariant under the exchange of $\sigma_{i}^{x}\leftrightarrow\tau_{i}^{z}$, $\sigma_{i}^{z}\leftrightarrow\tau_{i}^{x}$, and $h_{x}\leftrightarrow h_{z}$. Furthermore, its phase diagram is also symmetric about $h_{z}=h_{w}$, because $UH_{\text{dual}}(h_{x},h_{z},h_{w})U^{\dagger}=H_{\text{dual}}(h_{x},h_{w},h_{z})$, where the unitary transformation is $U=\prod_{i}\left(\frac{1+\sigma_{i}^{x}}{2}+\tau_{i}^{x}\frac{1-\sigma_{i}^{x}}{2}\right)$. Since duality transformations preserve the structure of phase diagrams, we can directly deduce the phase diagram of the $\mathbb{Z}_{4}$ QD model $H_{\text{QD}}$ from that of $H_{\text{dual}}$.
The non-local condensation order parameters and the Wilson loop expectation values of the perturbed $\mathbb{Z}_{4}$ QD model can be mapped to the local order parameters and the membrane-like disorder parameters of the two-coupled quantum Ising model [59, 60] (see SM [42] and Tab. 1). This allows us to extract condensation order parameters and the area law coefficients directly from the ground states of the two-coupled Ising model, which can be approximated using $C_{4v}$-symmetric iPEPS [37, 36, 38]. Compared to other methods such as quantum Monte Carlo, iPEPS is more suitable to evaluate non-local order parameters, and the method of extracting area law coefficients from iPEPS is proposed in Ref. [61]. By comparing the numerical results of the condensation order parameters and the area law coefficients with Tab. 1, we obtain the phase diagrams for various $h_{x}$ planes (see SM [42]), one of which is shown in Fig. 3. It includes the $\mathbb{Z}_{4}$ QD, TC, DS, and trivial phases, and their phase boundaries are described by the 3D Ising* universality class describing 1-form symmetry-breaking transitions [62]. Next, we focus on the multi-critical point, especially the continuous TC-DS transition.
@@
Fig. 4a shows that the area law coefficients $\alpha(\pmb{e})$, $\alpha(\pmb{e}\pmb{m})$ and $\alpha(\pmb{m})$ are zero when $h_{w}=h_{z}<h_{c}\approx 1.490$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{n}$, $\pmb{e}^{n}\pmb{m}^{k}$ and $\pmb{m}^{k}$ ($n,k\in\{1,3\}$) are confined. Moreover, Fig. 4b shows that the condensation order parameters $\langle\pmb{e}^{2}\rangle$, $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{m}^{2}\rangle$ are zero when $h_{w}=h_{z}<h_{c}$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ condense. By comparing the numerical results in Figs. 4a and b with Tab. 1, we conclude the perturbed $\mathbb{Z}_{4}$ QD model is in the $\mathbb{Z}_{4}$ QD when $h_{w}=h_{z}<h_{c}$ and trivial phases when $h_{w}=h_{z}>h_{c}$.